Subject: GrimII
From: Mark Klaene
Submitted: Tue, 23 Dec 2003 15:59:13 -0700
Message number: 730
(previous: 729,
next: 731
up: Index)
GrimII is currently back on the observing floor, cold and powered up.
It is available for use with the following limitation.
After a week of repair and troubleshooting by Dale Sandford it was
determined that the lower right quadrant failure is a result of a
problem inside the dewar. This quadrant was not able to be repaired
at this time. Much more extensive and invasive work will have to be
scheduled in the future. The 3 remaining quadrants appear usable and
nearly identical to before shutdown performance. There is still an
ongoing problem that causes any quadrant to occasionally drop out. A
proposed fix to this will be implemented in early January. Until then
we believe this is an infrequent occurrence and the instrument will
be useful.
For information about the state of the 3 active quadrants contact
John Barentine at the site.
--
Mark Klaene
Deputy Site Manager voice: 505 437 6822
PO Box 59 fax: 505 434 5555
Sunspot NM 88349-0059 web: www.apo.nmsu.edu
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 730 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO